Gabrielle's age is three times Mikhail's age. The sum of their ages is 64. What is Gabrielle's age?
Answer by ankor@dixie-net.com(12695)

Gabrielle's age is three times Mikhail's age.
G = 3M
:
The sum of their ages is 64.
G + M = 64
:
Repalce G with 3M
3M + M = 64
4M = 64
M = 16
:
What is Gabrielle's age?
G = 3(16)
G = 48